Project

General

Profile

News

Configuration Database: New release of Config DB Python API

Added by Wilson, Antony over 10 years ago

An updated Config DB Python API is available from launchpad

This corresponds to the server code that is on my test server
rgma19.pp.rl.ac.uk

The documentation for the API can be found here

N.B. Not all of the method will work with the current production servers.

Executive Board: Meeting 16 Dec 2011 (3 comments)

Added by Tunnell, Christopher over 10 years ago

Agenda:
  • actions and minutes (Tunnell, EB)
  • matters arising from the TB, and RF review (TB minutes were distributed by Tim Hayler, enclosed)
  • matters arising from the video conference
  • status and plans of constructions SS, AFC, CC, RF power ...
  • status of paper(s) (Maurizio)
  • data being public (Colling)
  • collaboration meetings in 2012
  • MPB (8 March) and FAC (19 March) meetings in 2012
  • common fund (Ken)
  • A.O.B.

MICE: MICE News 12 December 2011

Added by Kyberd, Paul over 10 years ago

Data Volume

As of run 3426, MICE has now taken more than a terabyte of data during Step 1!

MOM update

So far we have very successfully data taking period.<p>

Preliminary calibration of the TOF0 and TOF1 detectors shows that after
the refurbishment of TOF1 we have a significant improvement in the
resolution this station. Now both TOF0 and TOF1 have resolution of ~50 ps.

Progress has been made toward the understanding of the electron/positron
different time-of-flight puzzle. It has been demonstrated that variation
of the number of particle trigger per spill introduces difference in the
measured time-of-flight even when we use identical settings of the
beamline channel.

Limited Restaurant Service at RAL

On Tuesday 13 and Wednesday 14 December the Restaurant will be open for pre-booked Christmas lunches only.
The R22 coffee lounge will be open as normal. On 13 and 14 December the evening meal service will run as normal.

Alternative services:
A brunch service will be available from R22 from 07:30 to 11:00 (on 13 and 14 December)
The R22 Deli bar will be open as normal serving baguettes, rolls, sandwiches, Panini, soup,
jacket potatoes and a take away pasta dish (on 8, 13 and 14 December)
The R1 Coffee Lounge will be open as normal from 08:30 to 15:30 and provides a range of snacks.
Soup and jacket potatoes will also be available.
The Electron Cafe will be open from 08:30 to 15:30 and provides a range of snacks, one hot meal
during the lunch-time and Costa Coffee.

MICE: AFC Progress

Added by Kyberd, Paul over 10 years ago

Progress on the AFC is continuing with the successful potting of the first coil.

MICE: November 7th

Added by Kyberd, Paul over 10 years ago

Shifts

A call for shifters during the period of the run 1st-16th December has already gone out.
We also need shift cover for a number of days during the start up period:
21,22 November and 28,29,30 November. Although there is risk of a shift with little
to do in these periods, it is also the time when a lot can happen and successful
shifts during the startup period are vital if the December run is to be productive.
Please send requests to and check the shift pages
http://www.mice.iit.edu/mico/shifts/ for slots which still need to be filled.

MAUS developments

MAUS now has a dedicated IRC room, for communicating with the developers
in real time.
Network: irc.freenode.net
Room: #maus
For more details, follow the "IRC Channel" link from the MAUS wiki at
http://micewww.pp.rl.ac.uk/projects/maus/wiki/

MICE CM31 (from Alain Blondel)

Collaboration meeting in Mississippi: many thanks to Lucien Cremaldi,
David Sanders and Don Summers for hospitality. Beautiful scenes in
Mississippi including a very local "catfish and chips" collaboration dinner.
Much progress was reported and interesting discussions took place.
Some highlights (see also the presentations!)

-- Schedule was reviewed and usual slide produced.
Step I run will take place 1-16 December 2011. We are looking for a
second MOM to double on Yordan and for shifters. Discussion on physics
aims for this run at the analysis meeting Thursday 10 November 2011,
15:00 GMT/UT.
EMR delivery at RAL is now May 2012 due to improvements in design to
guarantee full efficiency. Possibility of EMR run will depend on
Spectrometer solenoid installation in Hall, which will take priority.
Delivery of Focus Coil at RAL is now end of February
Delivery of first Spectrometer Solenoid at RAL is now June 2012
Limited slip in Step IV start of data taking (now December 2012).,
delivery of
No definite date for step V and VI but many people are working hard to
keep step VI by 2016.
Still no definite plan for Coupling coil construction but lots of
progress: First bobbin was delivered in LBNL end October and large Dewar
for tests has now arrived at Fermilab.

-- small working group under leadership of Mike Courthold (Victoria
Blackmore, Roy Preece, Mike Zisman) will summarize the magnetic
environment studies and produce a concise report exposing the main
results and drawing the practical conclusions in terms of running MICE,
local shielding and operational precautions -- responsibilities if needed.

-- a working group will assemble under Linda Coney and Yordan Karadhzov
to recast the MICE timing and clocks, in order to have a single but
flexible reference for target, DAQ, TOFs, detectors and RF.

The current MICE schedule is at http://www.mice.iit.edu/news/2011/MICE-schedule-2011-10-31.pptx

Executive Board: Meeting 7 Oct 2011 (1 comment)

Added by Tunnell, Christopher over 10 years ago

Agenda
  • Actions and minutes (Chris Tunnel)
  • Matters from TB (Andy Nichols) (enclosed minutes)
  • agenda of upcoming collaboration meeting (Dan Kaplan-- Alain)
  • analysis meetings (Alain, John)
  • Status of Papers (Maurizio Bonesini)
  • talks etc... (Vittorio Palladino)
  • MOMS 2012
  • common fund (Ken Long)
  • aob

Executive Board: Meeting 16 September 2011 (1 comment)

Added by Tunnell, Christopher over 10 years ago

Agenda
  • Actions and minutes (Chris Tunnel)
  • Matters from TB (Andy Nichols)
  • Running plans
  • Agenda of upcoming collaboration meeting
  • Status of Papers (Maurizio Bonesini)
  • Talks etc... (Vittorio Palladino)
  • MOMS 2012
  • common fund (Ken Long)
  • aob

Executive Board: Meeting 12 August 2011 (1 comment)

Added by Tunnell, Christopher over 10 years ago

Agenda:

  1. Minutes of last meeting and review of actions
  2. Calendar
  3. Matters arising from TB (Nichols) how are we doing with milestones?
  4. Run schedule conclusions (Alain)
  5. Matters arising from VC (all)
  6. MICE Paper I report (Maurizio)
  7. Speakers board (Vittorio)
  8. Agenda of meeting in October (Oxford, Mississippi)
  9. Common fund (Ken Long)
  10. Any comments on the MPB report? (attached)
(101-110/140)

Also available in: Atom